# Statistical Taylor Expansion111

### v1

@@ 3 Variance Arithmetic
Variance arithmetic implements statistical Taylor expansion. Besides assuming that the input random variables are independently Gaussian distributed, it contains additional assumptions.
3.1 Numerical Representation
Variance arithmetic uses a pair of 64-bit standard floating-point numbers to represent the value $x$ and the variance $(\delta x)^{2}$ of an imprecise value $x\pm\delta x$. The predicted deviation $\delta x$ by variance arithmetic is named as uncertainty.
The equivalent floating-point value of the least significant bit of significand of a floating-point value is defined as the least significant value  [^7] The “least significant value” and “unit in the last place” are different in concept., which can be abbreviated as $LSV$ for simplicity. When the uncertainty $\delta x$ is not specified initially:
[…]
For an integer value: If it is within the range of $[-2^{53}+1,+2^{53}-1]$, it has 0 as its uncertainty; Otherwise, it is first converted to a conventional floating-point value before converted to an imprecise value.
@@ 3.2 Bounding Factor and Bounding Leakage
Formula (2.12) provides statistical bounding $\tilde{x}\in(x-\sigma\delta x,x+\sigma\delta x)$. If $\sigma=\infty$, $\zeta(2n)=(2n-1)!!$!, which causes Formula (2.16) to diverge for $\log(x\pm\delta x)$, and $(x\pm\delta x)^{c},c\in R$. Thus, $\sigma$ has to be limited, and statistical bounding is a necessity. When $\sigma$ is limited, Formula (2.12) becomes Formula (3.1):
@@
The property of $\zeta(2n)$ when $\sigma^{2}\ll 2n$ is not sensitive to the input uncertainty distribution. If the Normal distribution $N(\tilde{z)}$ in Formula (2.12) is replaced by a uniform distribution between $[-\sigma,+\sigma]$, Formula (3.4) is similar to Formula (3.3):
@@
Formula (3.5) gives the probability $\epsilon$ of $\tilde{x}\not\in(-\sigma\delta x,+\sigma\delta x)$, in which $\xi(z)$ is the cumulative density function for Normal distribution [5]. $\epsilon$ is defined as bounding leakage. When $\sigma=5$, $\epsilon=5.73\times 10^{-7}$, which is small enough for most applications. $\sigma=5$ is also a golden standard to assert a negative result [3].
@@ 3.3 Low-Order Approximation
Formula (2.22) (2.25), (2.6), and (2.31) can be simplified as Formula (3.6), (3.7), (3.3), (3.3), respectively.
@@ 3.4 Convergence
figure The measured upper bound $\overline{\delta x}$ for $\ln{x\pm\delta x}$ for different $x$ as shown by the x-axis, using the y-axis on the left. Also shown are the corresponding uncertainty bias and uncertainty, using the y-axis on the right.
@@
figure The measured upper bound $\overline{\delta x}$ for $(1\pm\delta x)^{c}$ for different $c$ as shown by the x-axis, using the y-axis on the left. Also shown are the corresponding uncertainty bias and uncertainty, using the y-axis on the right. When $c$ is a natural number, $\delta x$ has no upper bound, and such results are omitted in the figure.
@@
figure The measured upper bound $\overline{\delta x}$ for $\sin(x\pm\delta x)$ for different $x$ as shown by the x-axis, using the y-axis on the left. Also shown are the corresponding uncertainty bias and uncertainty, using the y-axis on the right. The unit of x-axis is $\pi$.
@@
figure The measured upper bound $\overline{\delta x}$ for $1/(1\pm\delta x)$ for different $\delta y$ of Taylor coefficient uncertainty as shown by the x-axis, using the y-axis on the left. Also shown are the corresponding uncertainty bias and uncertainty, using the y-axis on the right.
Due to digital limitation of conventional floating-point representation, $\zeta(2n)$ can only be calculated to $2442n<$ 244 before becoming infinitive, Thus, variance arithmetic requires different convergence rule from its counterpart of Taylor expansion of precise variable. In variance arithmetic, an infinitive expansion (including Taylor expansion) can fail for the following reasons: not monotonic, not stable, not reliable, and not finite.
@@
$\delta^{2}\ln(x)\simeq$ $\sum_{n=1}^{+\infty}P(x)^{2n}\zeta(2n)\sum_{j=1}^{2n-1}\frac{1}{j}\frac{1}{2n-j}=\sum_{n=1}^{+\infty}P(x)^{2n}\zeta(2n)\frac{1}{n}\sum_{j=1}^{2n-1}\frac{1}{j};$
$\lim_{n\rightarrow+\infty}P(x)^{2n}\zeta(2n)\frac{1}{n}\sum_{j=1}^{2n-1}\frac{1}{j}\simeq P(x)^{2n}\sigma^{2n}\frac{\ln(2n)}{(2n)^{2}}\sigma N(\sigma);$ (3.16)
@@
According to Formula (3.4), the condition for Formula (2.25) for $\log(x\pm\delta x)$ to converge is $P(x)<1/\sigma=1/5$, which is confirmed by Figure 3.4. Variance arithmetic rejects the distributional zero of $\ln(0)$ inside the range of $[x-\delta x,x+\delta x]$ statistically due to the divergence of Formula (2.25) mathematically, with $\zeta(2n)$ as the connection of these two worlds.
According to Formula (3.4), except when c is a nature number, Formula (2.31) for $(x\pm\delta x)^{c}$ converges near $P(x)\simeq 1/\sigma=1/5$, with the upper bound $\overline{P(x)}$ increasing with $c$. This is confirmed by Figure 3.4. However, due to complexity of Formula (2.31), there is yet any quantitative formula between largest $P(x)$ and $c$. Qualitatively, $\delta^{2}1/x$ converges slower than $\delta^{2}\sqrt{x}$ according to Formula (3.15) and (3.13).
Conventional condition to terminate an infinitive series numerically is when the residual of expansion is less than the $LSV$ of the result [13]. This method no longer works in variance arithmetic because the expansion is limited to $2442n<$ 244, and because the expansion result is no longer precise. Instead, to be sure that the expansion reaches its stable value statistically, if the absolute value of the last expansion term is more than $\tau$-fold of the result uncertainty, the result is invalidated as not stable, in which $\tau$ is a small probability related to $\sigma$ in Formula (3.18), and $\xi(z)$ is the cumulative density function for Normal distribution [5].
$2\xi(\tau)=\epsilon=2-2\xi(\frac{1}{2}+\sigma):\;\;\;\tau\simeq\epsilon\frac{\sqrt{2\pi}}{2}=7.18\times 10^{-7};$ (3.18)
Formula (2.30) shows that $\overline{x^{c}}$ converges slower with decreasing $c$. In Figure 3.4, beyond the upper bound $\overline{\delta x}$, the expansion is not stable when $c\leq-1.42$, or not monotonic when $c>-1.41$.
While $\delta x$ specifies the uncertainty of $x$, the calculated $\delta x$ itself can be imprecise. In this paper, to guarantee $\delta x$ to be precise enough, during expansion, if the value of the variance is less than the $1/\sigma$-fold of its uncertainty, the result is invalidated as not reliable.
The variance expansion can contain negative values, such as Formula (2.6) for $sin(x\pm\delta x)$. Figure 3.4 shows the upper bound $\overline{\delta x}$ for $sin(x\pm\delta x)$ is periodic between $0.318\pi$ and $0.416\pi$, which is nicely less than the periodicity $2\pi$ of $\sin(x)$. Beyond the largest possible $\delta x$, the expansion is invalidated as not reliable in Figure 3.4.
@@ 3.5 Continuity
When $c$ is a nature number $n$, Formula (2.31) has $n$ terms, so that $\delta x$ has no upper bound for $(x\pm\delta x)^{n}$, such as Formula (3.11) for $(x\pm\delta x)^{2}$. The result mean, variance and histogram of $(x\pm\delta x)^{c}$ are continuous around $c=n$.
@@
  • When the second derivative is negative, the resulted distribution is left-sided Delta-like, such as when $n=3,x=-0.2$. It is the mirror image of the distribution when $n=3,x=0.2$.
A statistical bounding range of variance arithmetic cannot include more than one distributional pole because the corresponding statistical Taylor expansion becomes not reliable. One such example is $\sin(x)$ in Figure 3.4.
@@ 3.6 Tracking Rounding Error
Variance arithmetic can track rounding errors effectively.
Figure 3.6 shows the residual error of $\sum_{j=0}^{121}x^{j}-\frac{1}{1-x}$, in which the polynomial $\sum_{j=0}^{121}x^{j}$ is calculated using Formula (2.37), $\frac{1}{1-x}$ is calculated using Formula (2.29), and $x$ is initiated as a floating-point value so that its uncertainty is the $LSV$ of $x$. Because $\eta(2n)$ is limited to $2n\leq 242$, Formula (2.37) is limited to $N\leq 121$, so that $\sum_{j=0}^{121}x^{j}$ has lower expansion order than that of $\frac{1}{1-x}$. Figure 3.6 shows:
  • When $x\in[-0.73,0.75]$, the required expansion order is no more than $121$, so that the residual error is caused by calculation rounding error between $\sum_{j=0}^{121}x^{j}$ and $\frac{1}{1-x}$. Detailed analysis shows that the max residual error is $4$-fold of the $LSV$ of $\frac{1}{1-x}$. The calculated uncertainty bounds the residual error nicely for all $x\in[-0.73,0.75]$.
  • When $x\not\in[-0.74,+0.75]$, such as when $x=-0.75,0.76$, the required expansion order is more than $121$, so that the residual error is caused by insufficient expansion order of $\sum_{j=0}^{121}x^{j}$, and the residual error increases when $|x|\rightarrow 1$.
@@ 3.7 Comparison
Two imprecise values can be compared statistically using their difference.
When the value difference is zero, the two imprecise values are equal. In statistics, such two imprecise values have $50\%$ possibility to be less than or greater to each other but no chance to be equal to each other [5]. In variance arithmetic, these two imprecise values are neither less nor greater than each other, so they are equal. In reality, the probability of two values equal exactly to each other is very low.
Otherwise, the standard z-stat method [5] is used to judge whether they are equal, or less or more than each other. For example, the difference between $1.002\pm 0.001$ and $1.000\pm 0.002$ is $0.002\pm 0.00224$, so that the z-value for the difference is $z=0.002/0.00224$, and the probability for them to be not equal is $2\xi(|z|)=62.8\%$, in which $\xi(z)$ is the cumulative density function for Normal distribution [5]. If the threshold probability of not equality is $50\%$, $1.000\pm 0.002<1.002\pm 0.001$. Instead of using the threshold probability, the equivalent bounding range for z can be used, such as $|z|\leq 0.67448975$ for the equal threshold probability of $50\%$.
@@ 4.1 Verification Methods and Standards
When an input contains unspecified errors, such as floating-point rounding errors, or numerical errors of library functions, Gaussian or white noises of increasing deviations can be added to the input, until ideal coverage is reached. The needed minimal noise deviation gives a good estimation of the amount of unspecified input uncertainties. The existence of ideal coverage is a necessary verification if the application of Formula (2.16) or (2.19) has been applied correctly for the application. The input noise range for the idea coverage defines the ideal application range of input uncertainties.
@@ 4.2 Types of Uncertainties
There are five sources of result uncertainty for a calculation [2][13]:
@@
Empirically, variance arithmetic can provide proper coverage to rounding errors, with the error deviations in the range of $[0.4,2.5]$.
Section 3 shows that variance arithmetic no longer needs theoretical reminder estimator [13]. Instead, it uses stability truncation, to guarantee that the truncation error is statistically small enough for the result uncertainty.
@@
The modelling errors arise when an approximate analytic solution is used, or when a real problem is simplified to obtain the solution. For example, Section 8 demonstrates that the discrete Fourier transformation is only an approximation for the mathematically defined Fourier transformation. Conceptually, modelling errors originate from mathematics, so they are outside the domain for variance arithmetic.

### v2

@@ 3 Variance Arithmetic
Variance arithmetic implements statistical Taylor expansion, by adding practical assumptions.
3.1 Gaussian Input Uncertainty Assumption
The central limit theorem converges other distributions toward Gaussian [5] after addition and subtraction, and such convergence is fast [1]. In a digital computer, multiplication is implemented as a series of shift and addition, division is implemented as a series of shift and subtraction, and general function is calculated as addition of expansion terms [8][9]. This is why in general an uncertainty without bounds is assumed to be Gaussian distributed [2][3][5]. Variance arithmetic also assumes that an imprecise input for statistical Taylor expansion is Gaussian distributed.
In variance arithmetic, $\zeta(2n+1)=0$.
@@ 3.2 Bounding Factor and Bounding Leakage
If $\sigma=\infty$, $\zeta(2n)=(2n-1)!!$!, which causes Formula (2.23) for $\log(x\pm\delta x)$ and Formula (2.26) for $(x\pm\delta x)^{c}$ to diverge. When $\sigma$ has an upper bound, Formula (2.7) becomes Formula (3.1):
@@
The property of $\zeta(2n)$ when $\sigma^{2}\ll 2n$ is not sensitive to the input uncertainty distribution. Formula (3.4) is the $\zeta(2n)$ for uniform distribution between $[-\sigma,+\sigma]$. It is similar to Formula (3.3):
@@
Formula (3.5) gives the probability $\epsilon$ of $\tilde{x}\not\in(-\sigma\delta x,+\sigma\delta x)$, in which $\xi(z)$ is the cumulative density function for Normal distribution [5]. $\epsilon$ is a probability defined as bounding leakage. When $\sigma=5$, $\epsilon=5.73\times 10^{-7}$, which is small enough for most applications. $\sigma=5$ is also a golden standard to assert a negative result [3].
@@ 3.3 Low-Order Approximation
When $\sigma=5,n<5$, Formula (2.20), and (2.23) can be simplified as Formula (3.6), and (3.7), respectively.
@@
To keep the result precision finer than $1$ for $e^{x\pm\delta x}$, $\delta x\lesssim 0.833$ regardless of $x$. In contrast, the result upper bound uncertainty $\delta\log(x\pm\delta x)\simeq 0.212$. Both are expected from the relation $\delta x\rightarrow P(e^{x})$, and $P(x)\rightarrow\delta\log(x)$, respectively.
When $\sigma=5,n<5$, Formula (2.26), and (2.29) can be simplified as Formula (3.3), (3.3), respectively.
$+\frac{7}{6}c^{2}(c-1)(c-2)^{2}(c-\frac{16}{7})P(x)^{6}+o(P(x)^{8});$ (3.8)
@@ 3.4 Convergence, Stability, Validity, and Reliability
figure The measured upper bound $\delta x$ for $\log{x\pm\delta x}$ for different $x$ as shown by the x-axis, using the y-axis on the left. Also shown are the corresponding uncertainty bias and uncertainty, using the y-axis on the right.
@@
figure The measured upper bound $\delta x$ for $(1\pm\delta x)^{c}$ for different $c$ as shown by the x-axis, using the y-axis on the left. Also shown are the corresponding uncertainty bias and uncertainty, using the y-axis on the right. When $c$ is a natural number, $\delta x$ has no upper bound, and such results are omitted in the figure.
@@
figure The measured upper bound $\delta x$ for $\sin(x\pm\delta x)$ for different $x$ as shown by the x-axis, using the y-axis on the left. Also shown are the corresponding uncertainty bias and uncertainty, using the y-axis on the right. The unit of x-axis is $\pi$.
@@
figure The measured upper bound $\delta x$ for $1/(1\pm\delta x)$ for different $\delta y$ of Taylor coefficient uncertainty as shown by the x-axis, using the y-axis on the left. Also shown are the corresponding uncertainty bias and uncertainty, using the y-axis on the right.
Due to digital limitation of conventional floating-point representation, $\zeta(2n)$ can only be calculated to $2442n<$ 244 before becoming infinitive, Thus, variance arithmetic has different convergence rule from its counterpart of Taylor expansion of precise variable. In variance arithmetic, an infinitive expansion (including Taylor expansion) can fail for the following reasons: not monotonic, not stable, not positive, not finite, and not reliable.
@@
$\delta^{2}\log(x)\simeq$ $\sum_{n=1}^{+\infty}P(x)^{2n}\zeta(2n)\sum_{j=1}^{2n-1}\frac{1}{j}\frac{1}{2n-j}=\sum_{n=1}^{+\infty}P(x)^{2n}\zeta(2n)\frac{1}{n}\sum_{j=1}^{2n-1}\frac{1}{j};$
$\lim_{n\rightarrow+\infty}P(x)^{2n}\zeta(2n)\frac{1}{n}\sum_{j=1}^{2n-1}\frac{1}{j}=P(x)^{2n}\sigma^{2n}\frac{\log(2n)}{(2n)^{2}}\sigma N(\sigma);$ (3.16)
@@
According to Formula (3.4), the condition for Formula (2.23) for $\log(x\pm\delta x)$ to converge is $P(x)\leq 1/\sigma=1/5$, which is confirmed by Figure 3.4. Variance arithmetic rejects the distributional zero of $\log(0)$ inside the range of $[x-\delta x,x+\delta x]$ statistically due to the divergence of Formula (2.23) mathematically, with $\zeta(2n)$ as the connection of these two worlds.
According to Formula (3.4), except when c is a nature number, Formula (2.26) for $(x\pm\delta x)^{c}$ converges near $P(x)\simeq 1/\sigma=1/5$, with the upper bound $P(x)$ increasing with $c$. This is confirmed by Figure 3.4. However, due to complexity of Formula (2.26), there is yet any quantitative formula between $\overline{P(x)}$ and $c$. Qualitatively, $\delta^{2}1/x$ converges slower than $\delta^{2}\sqrt{x}$ according to Formula (3.15) and (3.13).
Conventional condition to terminate an infinitive series numerically is when the residual of an expansion is less than the unit in the last place of the result [13]. In practice, the last expansion term usually proxies for the expansion residual if either the residual is known to be less than the last expansion term, or if the residual is difficult to calculate [13]. This method no longer works in variance arithmetic because the expansion is limited to $2442n<$ 244, and because the expansion result is no longer precise. The probability of an imprecie value $x\pm\delta x$ not equal 00 is indicated by its z-statistic $\frac{|x|}{\delta x}$ [5]. If this probability is the bounding leakage $\epsilon$, Formula (3.18) gives the corresponding upper bound $\tau$ for this z-statistic. Because variance arithmetic has already guaranteed that the expansion is absolutely and monotonically decreasing, if the absolute value of the last expansion term is more than $\tau$-fold of the result uncertainty, the result is invalidated as not stable.
$2\xi(\tau)=\epsilon:\;\;\;\tau\simeq\epsilon\frac{\sqrt{2\pi}}{2}=7.18\times 10^{-7};$ (3.18)
Formula (2.25) shows that $\overline{x^{c}}$ converges slower with decreasing $c$. In Figure 3.4, beyond the upper bound $\delta x$, the expansion is not stable when $c\leq-1.42$, or not monotonic when $c>-1.41$.
At any order, if the expansion variance becomes negative, the result is invalidated as not positive.
The variance expansion can contain negative values, such as Formula (2.29) for $sin(x\pm\delta x)$. Figure 3.4 shows that the upper bound $\delta x$ for $sin(x\pm\delta x)$ is periodic between $0.318\pi$ and $0.416\pi$, which is nicely less than the periodicity $2\pi$ of $\sin(x)$. Beyond the largest possible $\delta x$, the expansion is invalidated as not positive in Figure 3.4.
@@
While $\delta x$ specifies the uncertainty of $x$, the calculated $\delta x$ itself can be imprecise. To guarantee $\delta x$ to be precise enough, during expansion, if the value of the variance is less than the $1/\sigma$-fold of its uncertainty, the result is invalidated as not reliable.
Fortunately, being not reliable is rarely the reason to bound $\delta x$.
@@ 3.5 Continuity
The result mean, variance and histogram of variance arithmetic are generally continuous in parameter space.
When $c$ is a nature number $n$, Formula (2.26) has $n$ terms, so that $\delta x$ has no upper bound for $(x\pm\delta x)^{n}$ unless for not finite, such as Formula (3.11) for $(x\pm\delta x)^{2}$. The result mean, variance and histogram of $(x\pm\delta x)^{c}$ are continuous around $c=n$.
@@
  • When the second derivative is negative, the resulted distribution is left-sided Delta-like, such as when $n=3,x=-0.2$, which is the mirror image of the distribution when $n=3,x=0.2$.
  • In each case, the transition from $x=0$ to $x=0.2$ is continuous.
A statistical bounding range of variance arithmetic cannot include more than one distributional pole because the corresponding statistical Taylor expansion becomes not positive. One such example is $\sin(x)$ in Figure 3.4.
@@
3.6 Numerical Representation
Variance arithmetic uses a pair of 64-bit standard floating-point numbers to represent the value $x$ and the variance $(\delta x)^{2}$ of an imprecise value $x\pm\delta x$.
The equivalent floating-point value of the least significant bit of significand of a floating-point value is defined as the least significant value  [^6] The “least significant value” and “unit in the last place” are different in concept., which can be abbreviated as $LSV$ for simplicity.
When the uncertainty $\delta x$ is not specified initially:
[…]
  • For an integer value: If it is within the range of $[-2^{53}+1,+2^{53}-1]$  [^7] Including a hidden bit, the significand in the standard IEEE 64-bit floating-point representation is 53-bit., it has 0 as its uncertainty; Otherwise, it is first converted to a conventional floating-point value before converted to an imprecise value.
@@ 3.7 Tracking Rounding Error
Variance arithmetic can track rounding errors effectively without requiring additional rules.
Figure 3.7 shows the residual error of $\sum_{j=0}^{121}x^{j}-\frac{1}{1-x}$, in which the polynomial $\sum_{j=0}^{121}x^{j}$ is calculated using Formula (2.35), $\frac{1}{1-x}$ is calculated using Formula (2.24), and $x$ is initiated as a floating-point value. Because $\eta(2n)$ is limited to $2n\leq 242$, Formula (2.35) for polynomial is limited to $N\leq 121$, so that $\sum_{j=0}^{121}x^{j}$ has lower expansion order than that of $\frac{1}{1-x}$. Figure 3.7 shows:
  • When $x\in[-0.73,0.75]$, the required expansion order is no more than $121$, so that the residual error is the calculation rounding error between $\sum_{j=0}^{121}x^{j}$ and $\frac{1}{1-x}$. Detailed analysis shows that the max residual error is $4$-fold of the $LSV$ of $\frac{1}{1-x}$. The calculated uncertainty bounds the residual error nicely for all $x\in[-0.73,0.75]$.
  • When $x\not\in[-0.74,+0.75]$, such as when $x=-0.75,0.76$, the required expansion order is more than $121$, so that the residual error is caused by insufficient expansion order of $\sum_{j=0}^{121}x^{j}$, and the residual error increases absolutely when $|x|\rightarrow 1$, e.g., reaching $50$ when $x=0.98$.
@@ 3.8 Comparison
Two imprecise values can be compared statistically using the z-statistic [5] of their difference.
When the value difference is zero, the two imprecise values are equal. In statistics, such two imprecise values have $50\%$ possibility to be less than or greater to each other but no chance to be equal to each other [5]. In variance arithmetic, these two imprecise values are neither less nor greater than each other, so they are equal. In reality, the probability of two imprecise values equal exactly to each other is very low.
Otherwise, the standard z-statistic method [5] is used to judge whether they are equal, or less or more than each other. For example, the difference between $1.002\pm 0.001$ and $1.000\pm 0.002$ is $0.002\pm 0.00224$, so that the z-statistic for the difference is $z=0.002/0.00224$, and the probability for them to be not equal is $2\xi(|z|)=62.8\%$, in which $\xi(z)$ is the cumulative density function for Normal distribution [5]. If the threshold probability of not equality is $50\%$, $1.000\pm 0.002<1.002\pm 0.001$. Instead of using the threshold probability, the equivalent bounding range for z can be used, such as $|z|\leq 0.67448975$ for the equal threshold probability of $50\%$.
@@ 4.1 Verification Methods and Standards
When an input contains unspecified errors, such as numerical errors of library functions, Gaussian or white noises of increasing deviations can be added to the input, until ideal coverage is reached. The needed minimal noise deviation gives a good estimation of the amount of unspecified input uncertainties. The existence of ideal coverage is a necessary verification if the application of Formula (2.3) or (2.3) has been applied correctly for the application. The input noise range for the idea coverage defines the ideal application range of input uncertainties.
@@ 4.2 Types of Uncertainties
There are five sources of result uncertainty for a calculation [1][2][13]:
@@
Empirically, variance arithmetic can provide proper coverage to rounding errors, with the error deviations in the range of $[1/5,5]$.
Section 3 shows that variance arithmetic uses $\tau$-based rule for truncation error.
@@
The modelling errors arise when an approximate analytic solution is used, or when a real problem is simplified to obtain the solution. For example, Section 8 demonstrates that the discrete Fourier transformation is only an approximation for the mathematically defined continuous Fourier transformation. Conceptually, modeling errors originate from mathematics, so they are outside the domain for variance arithmetic.